two vectors of magnitute 20cm and 10cm make angles of 120 and 50 with Xaxis inthe XY PLANE FIND THEIR CROSS PRODUCT
1
Expert's answer
2020-09-01T10:37:40-0400
"\vec{a}\times\vec{b}=\begin{vmatrix}
 i & j & k \\
 20\cos{120} & 20\sin{120}& 0 \\
 10\cos{50} & 10\sin{50}& 0 
\end{vmatrix}\\=((20\cos{120})(10\sin{50})-(20\sin{120})(10\cos{50}))\vec{k}\\=(-188\vec{k})\ cm^2"